# Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis

### v3

@@ Abstract
Large Language Models (LLMs), already shown to ace various text comprehension tasks, have also remarkably been shown to tackle table comprehension tasks without specific training. Building on earlier studies of LLMs for tabular tasks, we probe how in-context learning (ICL), model scale, instruction tuning, and domain bias affect Tabular QA (TQA) robustness by testing LLMs, under diverse augmentations and perturbations, on diverse domains: Wikipedia-based WTQ, financial TAT-QA, and scientific SCITAB. Although instruction tuning and larger, newer LLMs deliver stronger, more robust TQA performance, data contamination and reliability issues, especially on WTQ, remain unresolved. Through an in-depth attention analysis, we reveal a strong correlation between perturbation-induced shifts in attention dispersion and the drops in performance, with sensitivity peaking in the model’s middle layers. We highlight the need for improved interpretable methodologies to develop more reliable LLMs for table comprehension.
Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis
@@
footnotetext: * These authors jointly supervised this work.
1 Introduction
LLMs, despite being primarily trained on unstructured text, have demonstrated notable capabilities in structured-data tasks such as Tabular Question Answering (TQA). TQA requires models to interpret data presented in tables, demanding strong structural reasoning. Assessing how LLMs navigate structured comprehension challenges can provide valuable insights into their robustness and reasoning capabilities Borisov et al. (2023); Fang et al. (2024).
Recent studies highlight the importance of robustness evaluations in understanding LLM behavior on structured tasks Zhou et al. (2024). Specifically, perturbations in tabular structure and content significantly impact model performance, highlighting vulnerabilities that limit practical reliability Wang et al. (2022); Zhao et al. (2023). Furthermore, Liu et al. (2023) argue that LLMs inherently struggle with structural manipulations, advocating for an integrated approach combining symbolic reasoning to enhance model robustness.
Figure 1: (a) An example of the different possible table augmentation methods.(b) Scatter plot between change in attention entropy and EM score across perturbation types for Llama-8B-Instruct on WTQ.
Although these studies identify vulnerabilities and suggest potential improvements, there remains limited understanding of how internal model mechanisms respond to perturbations in structured data. Attention mechanisms form the core of transformer-based LLMs, governing how models distribute focus across input elements during processing Clark et al. (2019). Prior work analyzing attention patterns in natural language tasks revealed that specific layers and attention heads critically influence model performance and robustness Zhao et al. (2024); Barbero et al. (2025). However, detailed attention-level analyses for structured tasks like TQA remain scarce.
To address this gap, we systematically investigate LLM robustness for TQA tasks across multiple dimensions, including in-context learning, model scale, instruction tuning, domain biases, value perturbations, and model size. Our experiments evaluate different perturbation types, highlighting how value-based alterations influence reasoning fidelity and faithfulness (Table 1). We also compare performance across diverse datasets to assess domain biases and generalizability.
Extending beyond surface-level performance metrics, we quantify changes in attention entropy across different attention heads and layers under various perturbations, analyzing how these changes correlate with performance degradation. With this we aim to uncover which attention heads are most sensitive to perturbations, informing targeted improvements for future models.
Our study leverages diverse datasets: Wikipedia-based WTQ Pasupat and Liang (2015), financial report-based TAT-QA Zhu et al. (2021), and scientific claims-based SCITAB Lu et al. (2023). Through these diverse domains, our findings give insights into the sensitivity and reliability of LLMs on TQA, highlighting the broader challenge of understanding how LLMs reason over structured data.
2 Various Perturbation Categories
Each perturbation is designed to manipulate table structure or content while preserving the inherent relational meaning of the table, to measure robustness to table comprehension, illustrated in Fig. 1. Structural Perturbation (SP): SP involves rearranging the columns and rows of the table to generate new examples. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap.
Value Perturbation (ValP): ValP focuses on modifying the actual data values within tables, ensuring that the model is faithful to the table. We explore these types of ValP:
Data Type Preserving Perturbation (DVP) [^1] Example of prompts and details of DVP dataset generation are present in the Appendix A, B: DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 5, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. These counterfactual entities test the faithfulness of the LLM to the tabular data. 222 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’.Random Value Perturbation (RVP): RVP (an example is shown in Fig. 6) relaxes DVP where instead of a counterfactual entity, we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’.
Null Value Perturbation (NVP): NVP removes the correct answer from the table completely. Evaluating the performance on NVP highlights the influence of Wikipedia content on solving the WTQ table question-answering task.
No Table (NT): To understand the extent of bias in WTQ, we evaluate LLMs on the no-table baseline. By analyzing the performance of LLMs in the absence of the table, we can better understand the extent of dependence on the particular tabular data.
@@ 3 Evaluation Performance
Operation WTQ Dataset TATQA Dataset SCITAB Dataset
Large Model Small Model Large Model Small Model Large Model Small Model
EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - - 0.33 - - 0.20 - - 0.098 - - 0.132 - -
Column 0.35 0.16 -0.01 0.25 0.15 0.00 0.31 0.09 -0.03 0.17 0.09 -0.03 0.073 0.085 -0.026 0.131 0.080 -0.001
Row 0.30 0.19 -0.06 0.20 0.17 -0.05 0.30 0.11 -0.03 0.18 0.09 -0.03 0.082 0.080 -0.017 0.132 0.079 0.000
Transpose 0.32 0.19 -0.04 0.20 0.18 -0.05 0.31 0.09 -0.02 0.18 0.09 -0.02 0.085 0.086 -0.015 0.134 0.082 0.002
Col Transpose 0.24 0.23 -0.12 0.15 0.20 -0.10 0.29 0.11 -0.05 0.16 0.10 -0.04 0.085 0.084 -0.013 0.130 0.082 -0.001
Row Transpose 0.31 0.20 -0.05 0.20 0.18 -0.05 0.26 0.13 -0.07 0.14 0.11 -0.06 0.088 0.085 -0.009 0.132 0.082 0.001
NT 0.06 0.36 -0.31 0.02 0.25 -0.23 0.01 0.33 -0.32 0.00 0.20 -0.20 0.000 0.113 -0.098 0.000 0.132 -0.132
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11 - - - - - - - - - - - -
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15 - - - - - - - - - - - -
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23 - - - - - - - - - - - -
Table 1: The average EM, VP, and Emd of small and large LLMs on all perturbation operations. Large and Small LLMs are defined on Table 3.
@@ 3.1 Effects of the Model Type on TQA
How do performances vary with value perturbations?  [^3] More details are in Appendices F and G Table 1 shows the performance of various LLMs, as defined on Table 3, when subjected to different value-based perturbation on WTQ using different evaluation metrics: Exact Match(EM), Variation percentage (VP) Yang et al. (2022); Zhou et al. (2024), and Exact Match Difference (EMD)Zhao et al. (2023); Zhou et al. (2024). We see LLMs experience a decline in EM scores across different operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency helps. VP increases significantly across all perturbations indicating considerable changes in predictions due to the perturbations. EMD consistently shows negative values, with the most substantial performance drops occurring in the NT and NVP scenarios. The observed discrepancies in performance, particularly pronounced in DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with perturbed tables.
@@ 3.2 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables? Table 1 shows that when no table is provided, LLMs show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5of queries correctly, indicating a potential bias in Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the models are biased by the contextual structured format cue even in the absence of relevant data.
How do out-of-box LLMs perform on specialized domains: TAT-QA and SCITAB ? As shown in Table 1, LLMs exhibit moderate performance across various table augmentations on the TAT-QA dataset and SCITAB dataset. For TAT-QA, large models consistently outperform smaller models, underscoring their superior TQA abilities, but for SCITAB, the small models have better performance in comparison to larger models. The overall inconsistent scores still denote significant challenges inherent in niche domains, but instruction-tuned models are better here. From Table 1, the comparison among WTQ, TAT-QA, and SCITAB datasets reveals an interesting accuracy-robustness tradeoff: while models with higher EM on WTQ suffer larger EMD and higher VP, indicating higher sensitivity to perturbations, their relatively smaller EMD and lower VP on both TAT-QA and SCITAB reflect greater robustness, despite SCITAB ’s overall lower baseline EM. Also notable is the sharp contrast in the NT performance, where on TAT-QA and SCITAB dataset, models have approximately 0% accuracy, a contrast from the WTQ performance. This suggests that performance on WTQ might be inflated due to biases favoring familiarity with Wikipedia, compared to niche domains like TAT-QA and SCITAB. This highlights the need for better benchmark design for tabular understanding. We also observe trends consistent with prior work: performance improves with few-shot prompting, instruction-tuned models generally outperform their base counterparts, and larger models tend to exhibit stronger TQA Wei et al. (2022); Fang et al. (2024). Additional results and full breakdowns across datasets, model sizes, and fine-tuning configurations are included in Appendix G.

### v4

@@ Abstract
Large Language Models (LLMs), already shown to ace various unstructured text comprehension tasks, have also remarkably been shown to tackle table (structured) comprehension tasks without specific training. Building on earlier studies of LLMs for tabular tasks, we probe how in-context learning (ICL), model scale, instruction tuning, and domain bias affect Tabular QA (TQA) robustness by testing LLMs, under diverse augmentations and perturbations, on diverse domains: Wikipedia-based WTQ, financial TAT-QA, and scientific SCITAB. Although instruction tuning and larger, newer LLMs deliver stronger, more robust TQA performance, data contamination and reliability issues, especially on WTQ, remain unresolved. Through an in-depth attention analysis, we reveal a strong correlation between perturbation-induced shifts in attention dispersion and the drops in performance, with sensitivity peaking in the model’s middle layers. We highlight the need for improved interpretable methodologies to develop more reliable LLMs for table comprehension. Through an in-depth attention analysis, we reveal a strong correlation between perturbation-induced shifts in attention dispersion and performance drops, with sensitivity peaking in the model’s middle layers. Based on these findings, we argue for the development of structure-aware self-attention mechanisms and domain-adaptive processing techniques to improve the transparency, generalization, and real-world reliability of LLMs on tabular data.
1 Introduction
LLMs, despite being primarily trained on unstructured text, have demonstrated notable capabilities in structured data tasks, such as Tabular Question Answering (TQA). TQA requires models to interpret data presented in tables, demanding strong structural reasoning. TQA specifically challenges models to discern relationships and hierarchies implicit within tabular data, making it an ideal benchmark for structural reasoning capabilities. Assessing how LLMs navigate structured comprehension challenges can provide valuable insights into their robustness and reasoning capabilities (Borisov et al., 2023; Fang et al., 2024). Tabular QA offers clean, controllable structural perturbations (e.g., row/column swaps, transposes) that directly target schema alignment and spatial reasoning while holding topical content constant. In contrast, long-context QA evaluations introduce confounds from retrieval, chunking, and position bias that make it harder to attribute attention-pattern changes to structure alone. Our focus on TQA thus isolates structure-sensitive behavior in a way that complements long-context frameworks such as NoLiMa (Modarressi et al., 2025).
Recent studies emphasize the importance of robustness evaluations in understanding LLM behavior on structured tasks (Zhou et al., 2024). Specifically, perturbations in tabular structure and content significantly impact model performance, highlighting vulnerabilities that limit the practical reliability of models (Wang et al., 2022; Zhao et al., 2023). Furthermore, Liu et al. (2023) argues that LLMs inherently struggle with structural manipulations, advocating for an integrated approach combining symbolic reasoning to enhance model robustness.
Figure 1: (a) An example of the different possible table augmentation methods.(b) Scatter plot between change in attention entropy and EM score across perturbation types for Llama-8B-Instruct on WTQ. The average attention entropy change against EM scores across perturbation types for Llama-8B-Instruct on WTQ, averaged over all attention heads and data points.
Although these studies identify vulnerabilities and suggest potential improvements, there remains limited understanding of how internal model mechanisms respond to perturbations in structured data. Attention mechanisms form the core of transformer-based LLMs, governing how models distribute focus across input elements during processing (Clark et al., 2019). Prior work analyzing attention patterns in natural language tasks revealed that specific layers and attention heads critically influence model performance and robustness (Zhao et al., 2024; Barbero et al., 2025). These attention heads often serve specialized functions such as syntactic parsing or semantic alignment, highlighting the complexity of internal transformer mechanisms. However, detailed attention-level analyses for structured tasks, such as TQA, remain scarce.
To address this gap, we systematically investigate LLM robustness for TQA tasks across multiple dimensions, including in-context learning, model scale, instruction tuning, domain biases, value perturbations, and model size. Our experiments evaluate different perturbation types, highlighting how value-based alterations influence reasoning fidelity and faithfulness (Table 1). We also compare performance across diverse datasets to assess domain biases and generalizability (detailed in Appendix A). Because robustness can be confounded by context length, we explicitly control for table size by restricting all inputs to <150<150 cells (Appendix D). We also group-report performance and attention statistics across table-size bins to probe scaling effects in a controlled manner (Appendix F; Fig. 13). Together, these controls allow us to clarify differences primarily to structural understanding rather than to generic long-context limitations (Modarressi et al., 2025).
Extending beyond surface-level performance metrics, we quantify changes in attention entropy across different attention heads and layers under various perturbations, analyzing how these changes correlate with performance degradation. Attention entropy effectively quantifies the dispersion in a model’s attention distribution, providing a nuanced metric for understanding internal decision-making processes. With this, we aim to determine which attention heads are most sensitive to perturbations, leading to more substantial changes, and thus more distinct performance degradation.
Our study leverages diverse datasets: Wikipedia-based WTQ (Pasupat & Liang, 2015), financial report-based TAT-QA (Zhu et al., 2021), and scientific claims-based SCITAB (Lu et al., 2023). Each of these datasets provides a unique context and complexity, allowing us to assess the generalizability and domain-specific sensitivities of LLMs robustly. Through these diverse domains, our findings give insights into the sensitivity and reliability of LLMs on TQA, highlighting the broader challenge of understanding how LLMs reason over structured
2 Various Perturbation Categories
Each perturbation is designed to manipulate the table structure or content while preserving the inherent relational meaning of the table, thereby measuring robustness to table comprehension, as illustrated in Fig. 1.
2.1 Structural Perturbation (SP):
SP involves rearranging the columns and rows of the table to generate new examples. These perturbations simulate realistic scenarios where data presentation varies significantly, testing the model’s ability to comprehend tabular structure. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap. These provide diverse perspectives on table comprehension, allowing for a thorough evaluation of the LLMs’ ability to handle structural variations.
2.2 Value Perturbation (ValP):
ValP focuses on modifying the actual data values within tables, ensuring that the model accurately reflects the data. Value perturbations specifically challenge the semantic fidelity of the model by altering critical data points that directly affect the answer. We explore these types of ValP:
  1. Data Type Preserving Perturbation (DVP): DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 9, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. These counterfactual entities test the faithfulness of the LLM to the tabular data. 111 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’. We utilize an automated counterfactual answer generation method that prompts GPT-3.5, ensuring the type correctness of the altered answer. Using a large language model, such fake answer generation makes it possible to generate fake answers that adhere to the data type and make the table semantically correct. Examples of prompts and details of DVP dataset generation are present in the Appendix B and C.
  2. Random Value Perturbation (RVP): RVP(an example is shown in Fig. 10) relaxes DVP where instead of a counterfactual entity, we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’. Performance on this perturbation correlates with whether the injection of random data into the table affects the accuracy. The comparison of random and data type-preserving perturbation also highlights whether models are influenced by the injection of abstract values for table comprehension.
  3. Null Value Perturbation (NVP): NVP removes the correct answer from the table completely. Evaluating the performance on NVP highlights the influence of Wikipedia content on solving the WTQ table question-answering task. LLMs that struggle more with the null value perturbation are likely to show consistent performance on TQA across different tabular datasets.
  4. No Table (NT): To understand the extent of bias in WTQ, we evaluate LLMs on the no-table baseline. This approach further emphasizes the reliance of LLMs on Wikipedia content. By analyzing the performance of LLMs in the absence of the table, we can better understand the extent of dependence on the particular tabular data. This method helps to reveal the intrinsic capabilities of LLMs for TQA and their generalizability across different tabular datasets.
To disentangle background memorization from table use, we employ four value-centric probes. NT and NVP bound performance when factual recall is necessary but local evidence is absent; DVP and RVP stress reliance on value-level fidelity when structure is intact. While a full label-randomization protocol would further isolate memorization effects, it introduces nontrivial semantic-consistency and schema-integrity issues in multi-column tables; we therefore leave a systematic randomization study to future work and use NT/NVP/DVP/RVP as practical, controlled approximations that foreground value-level dependency (related analyses in (Sakai et al., 2024)).
Collectively, these structural and value-based perturbations enable a comprehensive analysis of the model’s ability to reason over table structure and content. By introducing such constrained and adversarial transformations, we can more precisely isolate the underlying factors that drive the performance in TQA.
3 Evaluation Metrics
Given the definition of different perturbations, we employ these three metrics to evaluate model performance under various perturbations in TQA tasks.
  1. Exact Match Accuracy (EM): This metric calculates the proportion of instances for which the predicted answer exactly matches the ground truth answer. Formally, if $N$ is the total number of instances, and $\text{correct}(i)$ is an indicator function that is $1$ if the prediction for instance $i$ is correct and 0 otherwise, then:
    $\text{EM}=\frac{\sum_{i=1}^{N}\text{correct}(i)}{N}.$
  2. F1 Measure: The F1 measure is a harmonic mean of precision and recall, commonly used to evaluate the similarity between two text sequences, such as predicted and ground truth answers. Given two strings, each is tokenized into sets of words, and the overlap between them is used to compute:
@@ 3 Evaluation Metrics
    The F1 score is then computed as:
    $\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$
    A higher F1 score indicates greater similarity between the predicted and ground truth strings, while a lower F1 score reflects greater divergence.
  3. Exact Match Difference (Emd)(Zhao et al., 2023; Zhou et al., 2024): Let $\text{EM}_{\text{orig}}$ be the EM on the original (unperturbed) dataset, and $\text{EM}_{\text{perturbed}}$ be the EM after applying a perturbation. The Emd quantifies the change in EM due to perturbations:
    $\text{Emd}=\text{EM}_{\text{perturbed}}-\text{EM}_{\text{orig}}.$
  4. Variation Percentage (VP)(Yang et al., 2022; Zhou et al., 2024): This metric measures the extent to which predictions change after applying perturbations. Given that, C2W is the count of correct before perturbation and wrong after, and W2C is the count of wrong before perturbation and correct after. Given $N$ as the total number of instances, the variation percentage is:
    $\text{VP}=\frac{C2W+W2C}{N}.$
In addition to performance-based metrics, we also examine internal model behavior through the analysis of attention patterns.
  1. Attention Entropy: We analyze attention entropy to capture the dispersion of attention within different attention heads. Entropy serves as a measure for structural awareness, where high entropy indicates more evenly distributed attention, while low entropy reflects concentrated focus on a few tokens.
    $H_{i}=-\sum_{j}\textbf{A}_{ij}\log(\textbf{A}_{ij})$
    where $\textbf{A}_{ij}$ is the attention weight to the $j$-th token.
@@ 4 Evaluation Performance
Table 1: The average EM, F1, VP, and Emd of small and large LLMs on Value Perturbations(ValP). The classification of large and small models is defined in Table 3.
ValP Large Model Small Model
Operation EM F1 VP Emd EM F1 VP Emd
WTQ
Original 0.37 0.45 - - 0.26 0.34 - -
NT 0.05 0.06 0.37 -0.32 0.02 0.03 0.26 -0.24
DVP 0.22 0.29 0.40 -0.14 0.14 0.20 0.31 -0.12
RVP 0.15 0.20 0.39 -0.22 0.10 0.14 0.30 -0.16
NVP 0.06 0.09 0.37 -0.30 0.03 0.05 0.27 -0.23
TAT-QA
Original 0.34 0.43 - - 0.21 0.27 - -
NT 0.01 0.04 0.34 -0.33 0.00 0.02 0.20 -0.20
SCITAB
Original 0.113 0.116 - - 0.135 0.137 - -
NT 0.000 0.00 0.120 -0.120 0.000 0.00 0.132 -0.132
Figure 2: The average EM scores of the original/base(denoted by B) models and conversation(denoted by C)/instruction(denoted by I) models across various table augmentation techniques for WTQ, TAT-QA and SCITAB dataset with three fewshot examples.More individual results with additional metric are included in Appendix G.
4.1 Effects of ICL examples on TQA
Figure 3: Comparison of average EM scores across models under varying few-shot settings (0-shot, 1-shot, 2-shot, and 3-shot). Here, B denotes the Base variant of the model, C denotes the Chat variant, and I denotes the Instruction variant.
Does instruction prompting assist LLM for better table comprehension for question-answering tasks? The heat maps in Figure 2 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70B-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also yields better scores, albeit with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks. For Llama2 models, the gains from instruction-tuning are present but less pronounced compared to Llama3, especially on TAT-QA where the performance gap between base and instruction-tuned variants is narrower. Qwen2.5 models exhibit a similar improvement trend to Llama3, but their relative advantage is most visible on SCITAB, where instruction-tuned variants close a larger fraction of the base-to-top model gap. Qwen3 shows decent performance on WTQ, and also similar performance with the Base variant, suggesting that its base configuration already incorporates stronger reasoning priors for table tasks. It also distinctly indicates that LLMs that have undergone instruction or conversation-based fine-tuning outperform their base counterparts in TQA tasks for SCITAB dataset. Although SCITAB is inherently a classification task presented in a TQA format, using few-shot prompting provides valuable context, ultimately leading to more accurate and relevant responses.
@@ 4.2 Effects of the Model Type on TQA
Do newer models have better TQA abilities? Figure 4 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70B versions of these models generally perform better than their 7B variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities. While Llama3 models outperform their predecessors, the margin over Llama2 is largest on WTQ and narrows on TAT-QA and SCITAB. Notably, Qwen2.5 models, especially the larger variants, perform competitively with Llama3 on TAT-QA, often surpassing Mistral. For smaller variant models, Qwen3 shows comparable performance to Qwen2.5 model; however, the larger Qwen2.5 model retains an absolute performance edge in most cases.
Figure 4: The average EM for different models across the WTQ dataset with 3-fewshot settings over all structural perturbation. The models are categorized by their types (Llama2, Llama3, Mistral, Qwen2.5 and Qwen3) and separated by their sizes (7B, 70B, etc.)
Larger models (e.g., Llama3-70B, Mixtral-8x7B, Qwen2.5-72B) generally show higher performance than smaller models (e.g., Llama3-8B, Mistral-7B, Qwen2.5-8B), as seen in both the bar plot (Figure4) and heat maps (Figure2). For instance, Llama3-70B and Mixtral-8x7B have higher EM scores than Llama3-8B and Mistral-7B. However, we do see that average EM between the Qwen3 variant of small and large variants is not significantly different. This could be the result of an architectural difference between a dense model and MOE model. Although large, the active parameters in the 30B parameter model are only 3B parameter model. This suggests that model size has a significant impact on TQA performance.
How do performances vary with value perturbations?  [^2] More details are in Appendices F Table 1 shows the performance of various LLMs, as defined on Table 3, when subjected to different value-based perturbations on WTQ using different evaluation metrics: Exact Match(EM), Variation percentage (VP), and Exact Match Difference (EMD). We observe that LLMs experience a decline in EM scores across various operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency helps. VP increases significantly across all perturbations, indicating considerable changes in predictions due to the perturbations. EMD consistently shows negative values, with the most substantial performance drops occurring in the NT and NVP scenarios. The observed discrepancies in performance, particularly pronounced in DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with perturbed tables.
@@ 4.3 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables? Table 1 shows that when no table is provided, LLMs show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5% of queries correctly, indicating a potential bias in Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the models are biased by the contextual structured format cue even in the absence of relevant data.
How do out-of-box LLMs perform on specialized domains: TAT-QA and SCITAB? As shown in Table 1, LLMs exhibit moderate performance across various table augmentations on the TAT-QA dataset and SCITAB dataset. For TAT-QA, large models consistently outperform smaller models, underscoring their superior TQA abilities, but for SCITAB, the small models have better performance in comparison to larger models. The overall inconsistent scores still denote significant challenges inherent in niche domains, but instruction-tuned models are better here. From Table 1, the comparison among WTQ, TAT-QA, and SCITAB datasets reveals an interesting accuracy-robustness tradeoff: while models with higher EM on WTQ suffer larger EMD and higher VP, indicating higher sensitivity to perturbations, their relatively smaller EMD and lower VP on both TAT-QA and SCITAB reflect greater robustness, despite SCITAB ’s overall lower baseline EM. Also notable is the sharp contrast in the NT performance, where on TAT-QA and SCITAB datasets, models have approximately 0% accuracy, a contrast from the WTQ performance. This suggests that performance on WTQ might be inflated due to biases favoring familiarity with Wikipedia, compared to niche domains like TAT-QA and SCITAB. This highlights the need for better benchmark design for tabular understanding. We also observe trends consistent with prior work: performance improves with few-shot prompting, instruction-tuned models generally outperform their base counterparts, and larger models tend to exhibit stronger TQA (Wei et al., 2022; Fang et al., 2024).